OpenStudy (anonymous):

Prove or disprove: The set {(x, y),(a, b) | 3a + b = x, y = 2x − a} is a function from R × R to R × R.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

so the rule is (x,y) ----> (a,b) where 3a + b = x, y = 2x − a

OpenStudy (anonymous):

Would you begin by showing that a,b are arbitrary elements in RxR thus since x = 3a +b and y = 2(3a+b) - a then x,y must also be in RxR?

jimthompson5910 (jim_thompson5910):

so (a,b) is the input, (x,y) is the output?

jimthompson5910 (jim_thompson5910):

I'm not sure which is which

OpenStudy (anonymous):

Other way around, sorry. (a,b) is input

jimthompson5910 (jim_thompson5910):

so if this was a function, then (a,b) would map to EXACTLY ONE ordered pair (x,y)

OpenStudy (anonymous):

correct

jimthompson5910 (jim_thompson5910):

if (a,b) produced more than one ordered pair, then you can use that to disprove it's a function

OpenStudy (anonymous):

Correct, I cannot find any (a,b) that produces more than a single ordered pair, so how would I being proving this to be true?

jimthompson5910 (jim_thompson5910):

let's focus on 3a + b = x first

jimthompson5910 (jim_thompson5910):

3a + b is linear for a and b, so if you imagine setting up a 3-space axis (x, y, z axis), then 3a+b would be a plane in this 3-space setting

jimthompson5910 (jim_thompson5910):

for any given a and b, you get some height x, which is exactly one height I think

OpenStudy (anonymous):

correct

jimthompson5910 (jim_thompson5910):

so that leads me to think that 3a+b = x produces one value of x given any fixed (a,b) value

OpenStudy (anonymous):

Could you simply write (a,b) maps to (3a +b, 5a -b)?

OpenStudy (anonymous):

Could you simply write (a,b) maps to (3a +b, 5a + b)? rather

jimthompson5910 (jim_thompson5910):

how did you get the second value in the ordered pair?

jimthompson5910 (jim_thompson5910):

oh you plugged in x = 3a+b

jimthompson5910 (jim_thompson5910):

into y = 2x − a and simplified

OpenStudy (anonymous):

y = 2x - a, thus y = 2(3a +b) - a thus y = 6a + b - a, which means y = 5a + b

jimthompson5910 (jim_thompson5910):

i gotcha

jimthompson5910 (jim_thompson5910):

yeah a --> 3a+b which is basically mapping a line to a line b ---> 5a+b is also a line-to-line mapping

jimthompson5910 (jim_thompson5910):

so I think that's enough info to say that any given input produces one output

OpenStudy (anonymous):

and since they are linear there is only one input/output pair thus it must be a function by the horizontal line test?

jimthompson5910 (jim_thompson5910):

you mean vertical line test, but yes that's correct

OpenStudy (anonymous):

Yea sorry, my bad. Thank you for your help.

jimthompson5910 (jim_thompson5910):

you're welcome